### 19 Dec 2023 v1

# Cumulants and ordering of their ratios in 2D Potts models: Lessons for QCD?

Abstract
Theoretical considerations suggest an ordering of the ratios of net-baryon number fluctuations in the vicinity of the transition from the low-temperature hadronic phase to the high temperature quark-gluon plasma phase at small values of the baryon chemical potential, $\mu_{B}$, in the QCD phase diagram. The ordering hierarchy is $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, where $\chi_{n}$ is the $n^{\mathrm{th}}$ order cumulant of net-baryon number fluctuation. The STAR experiment observed this hierarchy in the ordering of cumulant ratios of net-proton number (a proxy of net-baryon number) for a range of colliding energies. These inequalities can be tested in spin models by taking the corresponding order parameters in the model as an analog of baryon density. We employed two different models: the two-state and three-state Potts models in two dimensions, which undergo a transition from an ordered phase to a disordered phase at their respective critical temperature. Simulations were performed on square lattices of different sizes using the Wolff algorithm. The cumulants of total magnetization are obtained up to the sixth order in both of these models in a temperature range near their corresponding critical temperatures. With increasing lattice size, height (trough) of the peaks (dips) of the higher-order cumulants appears to increase with the increase in the order of the cumulants. Except in a narrow range above the critical temperature of the three-state Potts model, the complete inequality or its complete reverse is not satisfied in the temperature ranges simulated.
I Introduction
Understanding the phase structure and the nature of phase transitions in the temperature-baryon density (or equivalently the baryon chemical potential, $\mu_{B}$) plane for the strongly interacting matter, known as quantum chromodynamics (QCD) phase diagram, is crucial for unraveling the fundamental properties of matter [1, 2]. While the QCD phase diagram is not well-explored for larger $\mu_{B}$, the region below $\mu_{B}$ = 450 MeV, spanning down to $\mu_{B}=0$, has been extensively studied, both in theory [2, 3] and in experiments [4]. For vanishing and small values of $\mu_{B}$, the transition from the low-temperature hadronic region to the quark-gluon plasma at high temperature is well-known to be a smooth transition (cross over) [5]. At larger values of $\mu_{B}$, however, it is generally expected that a first-order phase transition line exists, which ends in a second-order critical point belonging to the three-dimensional Ising model universality class [6, 7, 8]. Comprehensive studies of net-proton fluctuations have been carried out with the Beam Energy Scan (BES) program of the Relativistic Heavy Ion collider (RHIC) in the STAR experiment to search for the QCD critical point [9, 10]. Recently, an ordering of the cumulant ratios of net-proton number, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, has been reported by the STAR experiment across the energy range 7.7 to 200 GeV [11] where $\chi_{n}$ is the $n^{\mathrm{th}}$ order cumulant of event-by-event net-proton number distribution. This ordering was suggested previously by the Lattice-QCD (LQCD) calculations for the cumulants of net-baryon number fluctuations at small values of $\mu_{\mathrm{B}}$ [12, 13]. Functional renormalization group (FRG) calculations were also found to follow this hierarchy of the inequalities [14]. In contrast, predictions from the Hadron Resonance Gas (HRG) model, utilizing an ideal gas equation of state within a grand canonical ensemble framework, consistently yield positive unity for all cumulant ratios, lacking the observed ordering [15]. One may wonder whether these inequalities are generically valid in the vicinity of any phase transition or even in wider regions surrounding it. In the QCD phase diagram, either near the cross over or near the QCD critical point, if it exists, whether such inequalities exist is an interesting question to ask.
With the advent of lattice gauge theories, the Z(N) spin models have gained significant importance in the study of critical behavior in SU(N) gauge theories at finite temperature. Here, the variable N in general represents the number of degrees of freedom: within SU(N) gauge theories, it denotes the number of colors comprising the gauge group, while in Z(N) spin models, it represents the number of distinct spin states assignable to each lattice site. Among the Z(N) spin models, the three-state Potts model (Z(3)) in three dimensions (3D) holds particular significance as it provides a valuable framework for understanding the phase structures of QCD based on the SU(3) gauge symmetry group. The Polyakov loop, which serves as an order parameter of $\text{SU}(3)$ gauge theories, behaves similarly as spins in the low temperature phase of the 3D three-state Potts model [16, 17]. The line of first order phase transition in QCD phase diagram [18, 19] is closely related to the first order phase transition of 3D three-state Potts model in the absence of an external magnetic field [20, 17]. Addition of a small magnetic field ($h$) leads to a critical point in the Ising universality class for the three-state Potts model in 3D at $h_{c}$ [21], making it a nice toy model to explore the hierarchy of inequalities of the ratios of net-baryon number cumulants. As a first step towards such an investigation, we studied the higher-order cumulants of magnetization in the two-state (Z($2$)) and the three-state (Z($3$)) Potts model on a two-dimensional lattice in this paper. Both are known to have critical points albeit of another universality class. The finite-size scaling of the higher-order cumulants of magnetization were investigated near the critical region in the respective models and the ordering of their ratios were tested.
The paper is structured as follows: in Sec. II, we provide an overview of the $q$-state Potts models, detailing their formalism and the computation of cumulants of magnetization up to $6^{\mathrm{th}}$ order in these models. We also discuss the application of finite size scaling relation, which enables the extraction of exponents associated with the scaling of higher-order cumulants. Section  III presents the methodology employed for simulating the models and outlines the simulation validation process. In Sec. IV, we present the temperature-dependent behavior of cumulants and their ratios. The exponents involved with the finite-size scaling of the higher-order cumulants are estimated in terms of known critical exponents of the model and compared with the numerical results. Finally, Sec. V summarizes the key findings of our study.
II Model and Observables
The Hamiltonian for the the $q$-state Potts model [22] is given by
$\mathcal{H}=-\mathcal{J}\sum_{\langle i,j\rangle}\delta(\sigma_{i},\sigma_{j})-h\sum_{i}\delta(\sigma_{i},0)\,,$ (1)
[…]
$m_{\text{Potts}}=\frac{q}{q-1}\left(\frac{\text{max}(N_{0},N_{1},...,N_{q-1})}{N}-\frac{1}{q}\right)\,,$ (2)
where $N_{\alpha}=\sum_{j=1}^{L^{d}}\delta(\sigma_{i},\alpha)$ with $\alpha\in\{0,1,2,...,q-1\}$ represents the number of sites in the lattice with $\sigma_{i}=\alpha$ and $N=L^{d}=\sum_{\alpha}N_{\alpha}$ is the total number of spin sites in the lattice.
In the following analysis and results, we will be dealing Potts model with both $q$ = 2 and $q$ = 3 states. The two-state Potts model is equivalent to the well-known Ising model for which the exact solution by Onsager gives the critical point at $T_{C}=\frac{2}{\text{ln}(1+\sqrt{2})}\,\mathcal{J}/k_{\text{B}}\approx 2.2692\,\mathcal{J}/k_{\text{B}}$ [23]. The three-state Potts model that belongs to the same universality class as the two-dimensional Ising model, undergoes a second order phase transition at $T_{C}=\frac{1}{\text{ln}(1+\sqrt{3})}\,\mathcal{J}/k_{\text{B}}\approx 0.9950\,\mathcal{J}/k_{\text{B}}$ [22].
II.1 Cumulants
Figure 1: (a) The second order susceptibility of magnetization for the two-dimensional Ising model (equivalent to $q$ = 2 state Potts model) shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). (b) The reduced temperature, $t^{\prime}=T(\chi_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
In this paper, the main objects of interest are the higher-order cumulants (also referred as susceptibilities in the paper), which reflect the higher-order correlations amongst the spins. In spin models, the $p^{\text{th}}$ order susceptibility normalized by the total number of spin sites in the lattice is defined [24] by
$\chi_{p}(T)=\frac{1}{L^{d}\beta^{p}}\left(\frac{\partial^{p}\text{ln}Z}{\partial h^{p}}\right)_{h\to 0}\,,$ (3)
where $\beta=1/(k_{\text{B}}T)$, $T$ is the temperature, $k_{\text{B}}$ is the Boltzmann constant. Here, $Z$ is the partition function of the model given by its corresponding $\mathcal{H}$ as
$Z=\sum_{a}\text{exp}(-\beta\mathcal{H})\,.$ (4)
The sum over $a$ in Eq. 4 means the summation over all possible spin configurations on the lattice. The higher-order spin-spin susceptibilities can be associated with the cumulants of total magnetization ($M$). This is because $Z$ is proportional to the moment generating function [25] of $M$ and therfore one can easily deduce,
$\left(\frac{\partial^{p}\mathrm{ln}Z}{\partial h^{p}}\right)_{h\to 0}=\beta^{p}K_{p}(M)\,.$ (5)
Here, $M$ is referred as magnetization which is equal to the product of the order parameter and the volume of the lattice, i.e., $M=L^{d}m_{\mathrm{Potts}}$. $K_{p}(M)$ is the $p^{\text{th}}$ order cumulant of $M$. The expressions for $\chi_{p}$ (obtained by using the relation of cumulants in terms of central moments [26]) up to $p=6$ is listed below,
$\chi_{1}$ $=\langle M\rangle/L^{d}$ (6)
$\chi_{2}$ $=\mu_{2}/L^{d}$ (7)
$\chi_{3}$ $=\mu_{3}/L^{d}$ (8)
$\chi_{4}$ $=(\mu_{4}-3\mu_{2}^{2})/L^{d}$ (9)
$\chi_{5}$ $=(\mu_{5}-10\mu_{3}\mu_{2})/L^{d}$ (10)
$\chi_{6}$ $=(\mu_{6}-15\mu_{4}\mu_{2}-10\mu_{3}^{2}+30\mu_{2}^{3})/L^{d}$ (11)
where $\mu_{p}=\langle(M-\langle M\rangle)^{p}\rangle$ is the $p^{\text{th}}$ order central moment of the magnetization and the angular brackets, .$\langle$.$\rangle$ . . ⟩ represents average over all possible spin configurations on the lattice. We will be studying these $\chi_{p}$’s as a function of temperature in the critical region.
II.2 Finite size scaling
In this section, we obtain the generalized relation for finite size scaling of the higher-order susceptibilities of magnetization in the Potts model. We define, $t=(T-T_{C})/T_{C}$ and $h_{r}=\beta h$. For a second order (continuous) phase transition, there are at least two relevant scaling fields for the free energy: temperature like and magnetic field like. The dependence of free energy on these scaling fields near the critical point can be established by a bulk Renormalization Group (RG) transformation. RG transformation involves length scale re-scaling, under which, the scaling fields generally look like $C_{1}L^{1/\nu}t$ and $C_{2}L^{\Delta/\nu}h_{r}$ respectively [27, 28]. Here, the exponent $\nu$ is defined such that $\xi\sim t^{-\nu}$ (for a finite lattice $\xi\sim L$ at the critical point) and $\Delta\equiv\beta+\gamma$ is such that $\chi_{1}\sim t^{\beta}$ and $\chi_{2}\sim t^{\gamma}$ in the zero field limit, i.e., $h\rightarrow 0$.
In the obtained scaling relation for the singular part of the free energy after RG transformations, irrelevant scaling fields and non-linear contributions from relevant scaling fields can be safely removed by setting them to zero. This simplification is valid under the condition that $d<d_{>}$, where $d_{>}$ is the upper critical dimension. Consequently, only the linear order of temperature-like ($t$) and magnetic field-like ($h_{r}$) scaling fields remains, serving as the sole parameters that determine the distribution of the system [27]. The asymptotic finite size scaling relation of the singular part of reduced free energy density, i.e $f_{r}^{(s)}=L^{-d}\ln Z$, can therefore be written as (for $t\rightarrow 0,h_{r}\rightarrow 0$)
$f_{r}^{(s)}(t,h;L)\approx L^{-d}Y(C_{1}L^{1/\nu}t,C_{2}L^{\Delta/\nu}h_{r})\,.$ (12)
Here, the function $Y(x,y)$ is same for every system in the universality class, i.e. it is universal, while the coefficients $C_{1}$ and $C_{2}$ are generally not universal. The $n^{th}$ derivative of Eq. 12 with respect to $h_{r}$ gives the scaling relation for the $n^{th}$ order finite-size susceptibility,
$\chi_{n}^{(s)}(t;L)\approx-C_{2}^{n}L^{\frac{n\Delta}{\nu}-d}\frac{\partial^{n}Y}{\partial h_{r}^{n}}\bigg{|}_{h_{r}=0}\,.$ (13)
From Eq. 13, we can extract the scaling relation for the singular part of $\chi_{n}$, i.e., $\chi_{n}^{(s)}\sim L^{\frac{n\Delta}{\nu}-d}$. To simplify the exponents, we utilize the relations among the critical exponents: $\gamma=2\Delta-(2-\alpha)$ and $2-\alpha=d\nu$. Both of these relations are valid for $d=2$, but in general, the latter is only valid for any $d<d_{>}$. Thus, we can write the scaling exponent of $n^{\mathrm{th}}$ order susceptibility as,
$e_{n}\equiv\frac{n\Delta}{\nu}-d=\frac{1}{2}\frac{n\gamma}{\nu}+(n-2)d\,.$ (14)
These exponents for $n$ = 2 to 6 can now be calculated using the respective values of $\gamma$ and $\nu$ for the model under consideration. For the $q=2$ state Potts model on a two-dimensional lattice, we have $\gamma=\frac{7}{4}$ and $\nu=1$ for $q=2$ [29], while $\gamma=\frac{13}{9}$ and $\nu=\frac{5}{6}$ for $q=3$ [22].
Figure 2: (a) The second order susceptibility of magnetization for the two-dimensional $q$ = 3 state Potts model shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). (b) The reduced temperature, $t^{\prime}=T(C_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
III Numerical simulation
Both $q=2$ and $q=3$ state Potts model were simulated on two-dimensional lattices using the Wolff Cluster algorithm [30] at temperatures close to their corresponding critical temperature ($T_{C}$). In this algorithm, clusters of similarly oriented spin sites are determined randomly which are then flipped together to generate new states. This algorithm reduces the effect of critical slowing down near $T_{C}$ with increase in lattice size, i.e. it has a smaller dynamical exponent resulting in more efficient simulations in this region [31]. For $q=2$, square lattices of linear size $L$ = 50, 60, 70, 80 and 90 were used in the simulations and during the simulation $10^{4}$ independent configurations were generated after ensuring thermalization by throwing away sufficient initial configurations. Similarly, for $q=3$, simulations were performed on square lattices with $L$ = 48, 64, 80, 96 and 128 to generate $2\times 10^{4}$ independent configurations after approprite thermalization. In order to obtain independent spin configurations on the lattice, configurations were recorded at the interval of $2\tau$ where $\tau$ is the auto-correlation time that is obtained from the auto-correlation function (of $M$) defined in Eq. 15 using an exponential decay ansatz $\phi(t)\sim\mathrm{exp}\left({-\frac{t}{\tau}}\right)$.
$\phi(t)=\frac{\langle M(0)M(t)\rangle-\langle M\rangle^{2}}{\langle M^{2}\rangle-\langle M\rangle^{2}}$ (15)
The magnetization is calculated for each spin configuration in the simulation and the higher-order susceptibilities of magnetization are calculated using Eqs. 6 - 11 by averaging over all possible spin configurations on the lattice.
Figure 3: The third and fourth order susceptibilities of magnetization, $\chi_{3}$ and $\chi_{4}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to two-dimensional $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{3}$ and the bottom panel represents $\chi_{4}$ [where $\chi_{3}$ and $\chi_{4}$ divided by a factor $10^{5}$ ($10^{6}$) and $10^{8}$ ($10^{9}$) respectively for $q=2$ ($q=3$)].
IV Results
The top panels of Fig. 1 and Fig. 2 display the second-order susceptibility, denoted as $\chi_{2}$, for the $q=2$ and $q=3$ state Potts model respectively with varying lattice sizes, $L$. In both cases, $\chi_{2}$ exhibits a distinctive peak structure. As may be expected, the locations ($T_{C,L}$) and magnitudes ($\chi_{2}(T_{C,L})$) of these peaks vary with changes in the lattice size. Their scaling behavior can be utilized to extract the critical exponents, specifically $\gamma$ and $\nu$, which will be further elaborated upon in the subsequent section, IV.1.
Figure 4: The fifth and sixth order susceptibilities of magnetization, $\chi_{5}$ and $\chi_{6}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to two-dimensional $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{5}$ and the bottom panel represents $\chi_{6}$ [where $\chi_{5}$ and $\chi_{6}$ divided by a factor $10^{12}$ ($10^{13}$) and $10^{16}$ ($10^{17}$) respectively for $q=2$ ($q=3$)].
IV.1 Critical exponents
The critical exponents, $\gamma$ and $\nu$ can be obtained using the scaling relations, $\chi_{2}(T_{C,L})\propto L^{-\gamma/\nu}$ and $T_{C,L}-T_{C}\propto L^{-1/\nu}$, where $T_{C,L}$ is the psuedo-critical temperature for a given lattice size, and $T_{C}$ represents the critical temperature of the model in the limit $L\to\infty$. The pseudo-critical temperature is an effective critical temperature observed in finite-sized systems, reflecting behavior similar to the critical point in infinite-sized systems. $T_{C,L}$ is defined as the location of the peak in the second order susceptibility, $\chi_{2}$ distribution as shown in Figs. 1 and  2. By fitting $\chi_{2}$ as a Gaussian function of the temperature, $T$, the peak location ($T_{C,L}$, $\chi_{2}(T_{C,L})$) is determined for each lattice size. The obtained values of $T_{C,L}$ and $\chi_{2}(T_{C,L})$ as a function of $L$ are tabulated in Tables 1 and 2 for $q$ = 2 and $q$ = 3 state Potts model respectively. Also, we define the critical region around $T_{C,L}$ for each lattice size by the width of the $\chi_{2}$ distribution peak, which is equal to the $\sigma$ of the fitted Gaussian distribution function, represented here as $\sigma_{C}$. Using the values of $T_{C,L}$, the difference $t^{\prime}=T_{C,L}-T_{C}$ is plotted as a function of $1/L$, and fitted with its scaling relations for both $q=2$ and $q=3$ state Potts model as shown in the bottom panels of Figs. 1 and 2 respectively. Alternatively, $T_{C}$ can be estimated by plotting directly $T_{C,L}$ against $L$ and fitting with $y=A+Bx^{-1/\nu}$. The resulting estimates ($T_{C}=2.2780\pm 0.0089$ for $q=2$ and $T_{C}=0.9953\pm 0.0001$ for $q=3$) are found to be consistent with the known values [23, 22] within uncertainties. For precise measurements of the critical exponents, however we employ the known values of $T_{C}$ in the former method and extracted the coefficient $\nu$. The peak position, $\chi_{2}(T_{C,L})$ as a function of $L$, is also fitted with its scaling relation. Therefore, the extracted values of $\gamma$ and $\nu$ from the fits with the scaling relations are as follows: for $q=2$, $\gamma=1.758\pm 0.060$, $\nu=1.019\pm 0.034$ whereas for $q=3$, $\gamma=1.435\pm 0.014$, $\nu=0.830\pm 0.008$. The excellent agreement observed between the calculated values of $\gamma$ and $\nu$ and their theoretical counterparts affirms the validation of our simulation.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C}$
50 2.3151 $\pm$ 0.0007 105.656 0.0495 $\pm$ 0.0011
60 2.3076 $\pm$ 0.0006 143.913 0.0424 $\pm$ 0.0009
70 2.3020 $\pm$ 0.0007 188.497 0.0360 $\pm$ 0.0009
80 2.2974 $\pm$ 0.0005 238.766 0.0310 $\pm$ 0.0007
90 2.2956 $\pm$ 0.0005 289.739 0.0288 $\pm$ 0.0007
Table 1: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C}$) of the peak of $\chi_{2}$ distribution as a function of temperature for $q$ = 2 state Potts model. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C}$
48 1.0039 $\pm$ 0.0001 105.656 0.0102 $\pm$ 0.0002
64 1.0012 $\pm$ 0.0001 143.913 0.0071 $\pm$ 0.0002
80 0.9998 $\pm$ 0.0001 188.497 0.0056 $\pm$ 0.0001
96 0.9989 $\pm$ 0.0001 238.766 0.0046 $\pm$ 0.0001
128 0.9978 $\pm$ 0.0001 289.739 0.0033 $\pm$ 0.0001
Table 2: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C}$) of the peak of $\chi_{2}$ distribution as a function of temperature for $q$ = 3 state Potts model. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
We plot the higher-order susceptibilities, $\chi_{3}$ and $\chi_{4}$ for both $q=2$ and $q=3$ state Potts model in Fig. 3 as a function of temperature scaled by $T_{C,L}$. Since, these $\chi_{n}$’s are one order derivative of its preceding $\chi_{n-1}$ (for $n>2$), the sign of peak structure gets reversed in $\chi_{n}$ with respect to $\chi_{n-1}$. Similarly, $\chi_{5}$ and $\chi_{6}$ are shown as a function of $T/T_{C,L}$ for both the models in Fig. 4. The height of the peaks or dips in these higher-order susceptibilities are found to be lattice size dependent. They also follow a scaling as a function of $L$ as discussed in section II.2. In order to determine the scaling exponents of the higher-order susceptibilities, we fit the maximum or minimum value of $\chi_{n}$, depending on which is prominent, as a function of $L$ against the scaling relation: $\chi_{n}^{\text{max/min}}\propto L^{b}$. The scaling exponents of $\chi_{3}$, $\chi_{4}$, $\chi_{5}$ and $\chi_{6}$ obtained from the fit are referred as $e_{n}^{{}^{\prime\prime}}$. For $q=2$ state Potts model, the values of scaling exponents are given in Table 3, whereas for $q=3$ state Potts model, they are provided in Table 4. Additionally, we compare the scaling exponents (Eq. 14) obtained using exact values of $\gamma$ and $\nu$ as well as by using the estimates of $\gamma$ and $\nu$ (from scaling of $\chi_{2}$) obtained in this paper, and they are referred as $e_{n}$ and $e_{n}^{\prime}$ respectively in Tables 3 and 4. The numerically estimated scaling exponents, $e_{n}^{\prime\prime}$ for the higher-order susceptibilities show a good agreement with the analytically obtained values. We also note that determining the exponents from finite size scaling of $\chi_{n}$’s instead of using the numerical estimate of $\gamma$ and $\nu$ from $\chi_{2}$ result in smaller numerical errors.
Quantity $e_{n}$ $e_{n}^{\prime}$ $e_{n}^{\prime\prime}$
$\chi_{3}$ 3.625 $3.566\pm 0.097$ $3.608\pm 0.026$
$\chi_{4}$ 5.500 $5.420\pm 0.130$ $5.500\pm 0.074$
$\chi_{5}$ 7.375 $7.280\pm 0.160$ $7.257\pm 0.067$
$\chi_{6}$ 9.250 $9.130\pm 0.190$ $9.196\pm 0.066$
Table 3: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in two-dimensional $q$ = 2 state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=7/4$ and $\nu=1$) in Eq. 14, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. IV.1. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
Quantity $e_{n}$ $e_{n}^{\prime}$ $e_{n}^{\prime\prime}$
$\chi_{3}$ 3.599 $3.593\pm 0.036$ $3.619\pm 0.021$
$\chi_{4}$ 5.466 $5.458\pm 0.047$ $5.501\pm 0.047$
$\chi_{5}$ 7.333 $7.322\pm 0.059$ $7.301\pm 0.047$
$\chi_{6}$ 9.199 $9.187\pm 0.071$ $9.197\pm 0.077$
Table 4: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in two-dimensional $q$ = 3 state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=13/9$ and $\nu=5/6$) in Eq. 14, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. IV.1. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
Figure 5: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for two-dimensional $q$ = 2 state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded region represents the critical region, $T_{C,L}\pm\sigma_{C}$.
Figure 6: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for two-dimensional $q$ = 3 state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded region represents the critical region, $T_{C,L}\pm\sigma_{C}$.
IV.2 Ordering of Cumulant ratios
In this section, we delve into the discussion of whether the two-dimensional $q=2$ and $q=3$ state Potts models exhibit ordering in the susceptibility ratios, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$. Figure 5 presents a comparison of susceptibility ratios, $\frac{\chi_{3}}{\chi_{1}}$, $\frac{\chi_{4}}{\chi_{2}}$, $\frac{\chi_{5}}{\chi_{1}}$, and $\frac{\chi_{6}}{\chi_{2}}$, in the $q$ = 2 state Potts model for lattice sizes of 50, 60, and 80 as a function of temperature scaled by $T_{C,L}$. The shaded yellow band corresponds to the critical region, $T_{C,L}\pm\sigma_{C}$, where the values of $\sigma_{C}$ are listed in Table 1. It is already clear from Figs. 3 and 4 that the scale of the susceptibilities differ by factors larger than $\sim\mathcal{O}(10^{3})$. Therefore, to investigate the inequalities among the susceptibility ratios (if it exists) thoroughly, the ratios are displayed pairwise in three panels (for easy inspection by eye). Furthermore, a meaningful discussion of the ordering of the ratio of susceptibilities is feasible in the critical region around $T_{C,L}$ as a result of the vast changes in the magnitudes of successively higher orders. A clear ordering, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ is observed for $T>T_{a}$ in the critical region from the top and middle panel plots of Fig. 5, where $T_{a}\sim 1.01T_{C,L}$ ($T_{a}\sim 1.005T_{C,L}$) for $L=50$ ($L=80$). But the ordering, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}$ is not observed for $T>T_{a}$. Instead, a reverse ordering is observed, with $\frac{\chi_{5}}{\chi_{1}}$ being smaller than $\frac{\chi_{6}}{\chi_{2}}$ for the entire $T>T_{C,L}$. Therefore, we find that the expected complete ordering as seen in [11] is not observed in any temperature range near the critical region in $q=2$ state Potts model. Only a truncated ordering of the susceptibility ratios up to $\frac{\chi_{5}}{\chi_{1}}$, i.e., $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ holds true for $T>1.01T_{C,L}$ across all lattice sizes shown in Fig. 5. Such a truncated inequality is not observed for the entire temperature range, $T<T_{a}$.
[…]
A partial reversal of the ordering up to $\frac{\chi_{5}}{\chi_{1}}$, i.e., $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$ is seen in a narrow temperature range below the $T_{C,L}$, but a complete reversal is not observed for both $q=2$ and $q=3$ state Potts model. An important point to note is that regions of inequality exclusively lie in the critical region, which itself shrinks as in the infinite volume limit. This suggests that for infinite system size, we may not be able to establish any inequalities for the susceptibility ratios of magnetization in the spin models considered, and any observed orderings are a finite size effect. It would be interesting to check if any such volume dependence of the produced medium is seen in the relativistic heavy ion collisions.
V Conclusions
LQCD predicts a specific hierarchy of net-baryon number susceptibility ratios near the quark-hadron transition temperature for small $\mu_{B}$: $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$. The STAR experiment tested these predictions for net-proton number cumulant ratios in Au+Au collisions at various center-of-mass energies. Cumulant ratios from 7.7 to 200 GeV exhibited the predicted hierarchy, while a reverse ordering was observed at 3 GeV. If these inequalities are universally applicable for any system undergoing phase transition, is what we tried to understand.
In this study, we investigated higher-order susceptibilities of magnetization and their ratios in the critical regions of two-dimensional two and three-state Potts models. Utilizing the Wolff cluster algorithm to mitigate critical slowing down effects, simulations were conducted. We employed finite size scaling techniques to extract scaling exponents associated with higher-order susceptibilities, extending up to the sixth order. Additionally, we estimated these scaling exponents using the known critical exponents of the model, specifically the $\nu$ and $\gamma$ exponents. The obtained values from both methods exhibited a good agreement, indicating that the behavior of higher-order susceptibilities aligns well with our understanding of second-order phase transitions in the spin models.
We observed an absence of a consistent order in susceptibility ratios both above and below the pseudo-critical point (pseudo-critical temperature). A truncated ordering, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ is observed in a small temperature range, greater than the pseudo-critical temperature for a given lattice size in both two and three-state Potts models. In a yet narrow temperature range just above the pseudo-critical temperature, the complete ordering hierarchy in the susceptibility ratios is noted in the three-state Potts models for $L=128$ (illustrated in Fig. 6). While a complete reversal of this ordering is not observed in either of the models, we acknowledge that a partial reversal of the ordering, $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$ is seen in a small temperature range below the pseudo-critical temperature. Furthermore, we found that the region suitable for making comparisons between these ratios diminishes notably as the lattice size (or “lattice volume”) increases, shrinking to a point in the infinite volume limit. Consequently, the orderings observed in the susceptibility ratios can be attributed to finite size effects in spin model.
The spin models considered here have only a second order phase transition. A more realistic simulation which will be in better analogy with the expectations for the QCD phase diagram entails employing $q=3$ state Potts model on a three-dimensional lattice and with a small but non-vanishing magnetic field. It will be interesting to see what changes does a first order phase transition or the cross over region beyond the critical magnetic field make to the above observations on the inequality structure.
Acknowledgements.
We acknowledge the support of the Department of Atomic Energy (DAE) and the Department of Science and Technology (DST), Government of India. RVG expresses gratitude for the financial support from DAE through the Raja Ramanna Fellowship and the hospitality of the Department of Physics at IISER Bhopal. BM and SS acknowledge the financial support of DAE and DST. We also acknowledge the use of Garuda HPC facility at the School of Physical Sciences, NISER. SS would like to thank V K S Kashyap for technical help in running simulations, as well as Ashish Pandav and Debasish Mallick for insightful discussions regarding STAR results.